what is the slope point for the line given point that has been given
(5,8);m=-3

Answer by stanbon(75887)

what is the slope point for the line given point that has been given
(5,8);m=-3
------------------------
Form: y = mx + b
You are given m=-3
You are given y = 8 when x = 5
Substitute and solve for "b:
8 = -3*5 + b
8 = -15 + b
b = 23
============
Equation:
y = -3x + 23
============
Cheers,
Stan H.